OpenStudy (anonymous):

The sum of the interior angles of a hexagon is ____ 1,260° 1,080° 900° 720°

OpenStudy (anonymous):

are you sure those are the right choices?

OpenStudy (anonymous):

fixed

OpenStudy (welshfella):

we are talking angles here not lengths

OpenStudy (anonymous):

answer isnt there

OpenStudy (anonymous):

180(n-2) is the formula for the sum of interior angles of a polygon with n sides

OpenStudy (anonymous):

i hate this >.<

OpenStudy (welshfella):

the formula for the total angles in a polygon of side n is 180 ( n - 2)

OpenStudy (welshfella):

* n sided polygon

OpenStudy (welshfella):

so for a 4 sided polygon its 180(4 - 2) = 180*2 = 360 degrees

OpenStudy (welshfella):

a hexagon has 6 sides

OpenStudy (anonymous):

so its 180*4 for this one?

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

ooh i was thinking 180*40 for some reason lol

OpenStudy (anonymous):

so 720

OpenStudy (anonymous):

\(\checkmark\)

OpenStudy (loser66):

baby steps: 3 sides --- 180 every 1more side (180) + 180 hence 4sides --- 180+180 =360 5sides ---360+180 =540 6sides--- 540 +180= 720 ............................................

OpenStudy (anonymous):

:3